### v2

# Towards Fair In-Context Learning with Tabular Foundation Models

Abstract
Tabular foundational models have exhibited strong in-context learning (ICL) capabilities on structured data, allowing them to make accurate predictions on test sets without parameter updates, using training examples as context. This emerging approach positions itself as a competitive alternative to traditional gradient-boosted tree methods. However, while biases in conventional machine learning models are well documented, it remains unclear how these biases manifest in tabular ICL. The paper investigates the fairness implications of tabular ICL and explores three preprocessing strategies—correlation removal, group-balanced demonstration selection, and uncertainty-based demonstration selection—to address bias. Comprehensive experiments indicate that uncertainty-based demonstration selection consistently enhances group fairness of in-context predictions. The source code for reproducing the results of this work can be found at https://github.com/patrikken/Fair-TabICL.
1 Introduction
Tabular data, represented in rows and columns, is a data modality widely used for prediction tasks in domains such as finance and healthcare (Asuncion et al., 2007). Tree-based models such as XGboost (Chen et al., 2015) and Gradient-Boosted Trees (Ke et al., 2017) have shown the strongest generalization performance on tabular data. Recently, with the emergence of foundation models, Deep Learning (DL) based models have challenged the dominance of tree-based models (Hollmann et al., 2025). Foundation models are transformer models (Vaswani et al., 2017) pretrained on massive amounts of data, learning various complex structures that enable in-context learning (ICL) with a few labelled data (Brown et al., 2020). In-context learning has mainly been used with large language models (LLMs) for natural language tasks, where for a given prediction task, the labelled samples are formatted into textual demonstration examples and provided as context information to a language model such as GTP-3 (Radford et al., 2019). Given the demonstrations as context, the language model can effectively predict the label of a test example without any model update or finetuning (Brown et al., 2020). Attempts have been made to perform ICL using LLMs on tabular data, where the rows of the table are serialized into text or sentences (Hegselmann et al., 2023). However, LLMs are not pretrained to fully capture the complex relations between rows and columns of tabular data. Even when fine-tuned on vast amounts of tabular data, LLM-based ICL is limited by the size of the context window (up to 32 or 64 shots). To overcome this, Hollmann et al. (2022) proposed a TabPFN, a transformer-based tabular foundation model fully pretrained on synthetic datasets. Followup versions of TabPFN—TabPFNv2 (Hollmann et al., 2025) and TabICL (Qu et al., 2025)— demonstrated competitive or better performance compared to the traditional tree-based model on a variety of tasks under ICL (Hollmann et al., 2025; Qu et al., 2025).
With their state-of-the-art performance and ICL capabilities, transformer-based tabular foundation models will likely be widely adopted for decision-making and trigger a shift in the learning paradigm. However, using ICL-based models in high-stakes decision-making settings requires a thorough investigation of potential negative social impact, e.g., discrimination based on sensitive attributes such as gender, race, or sexual orientation. It has been shown that traditional machine-learning models can perpetuate bias in the data (Mehrabi et al., 2022) even when demographic features are not used to train the model (Kenfack et al., 2024a). Recent studies have demonstrated that ICL can also provide biased predictions (Hu et al., 2024; Bhaila et al., 2024). However, these works focus on LLM-based ICL with tabular data serialized into text format, thereby inheriting the drawbacks of foundational language models on tabular data (Hollmann et al., 2025). Moreover, tabular data is the main modality found in the fairness literature (Le Quy et al., 2022), and studies on fair ICL should use foundation models that can achieve state-of-the-art performance for tabular datasets (Qu et al., 2025).
This paper investigates the fairness of ICL prediction using transformer-based tabular foundation models. First, our study reveals, perhaps unsurprisingly, that while these models focus on improving prediction accuracy, they can also amplify bias. Motivated by recent studies on the sensitivity of ICL performance—in terms of fairness and accuracy— to demonstration selection, we aim to address the following research question: What in-context selection/transformation method can improve the fairness of ICL predictions? In the fairness literature, fairness-enhancing methods are generally grouped into three categories: pre-processing, in-processing, and post-processing (Mehrabi et al., 2022). Projecting these categories into the ICL paradigm, pre-processing methods perform demonstration transformation or selection before predicting in context (Hu et al., 2024). In-processing methods would fine-tune or retrain the foundation model with fairness constraints (Robertson et al., 2024). Post-processing methods would alter the ICL predictions to improve a given fairness metric (Hardt et al., 2016). Pre- and Post-processing methods are more computationally friendly since they do not require model updates. This motivates our choice to focus on the pre-processing techniques, and we leave post-processing interventions for future exploration. More specifically, we propose and investigate three pre-processing fairness interventions: (i) Correlation Remover, a method that alters each non-sensitive feature to reduce their correlation with the sensitive attribute; (ii) group-balanced111Underlined represent the method’s name throughout the paper and in the results. in-context selection, ensures that the in-context set is group-balanced; (iii) Uncertainty-based in-context selection, estimates the uncertainty of predicting the sensitive attribute of in-context samples and only selects samples with uncertain predictions. We provide intensive experiments on eight fairness benchmark datasets to investigate the effectiveness of each method in terms of fairness and accuracy. Our results reveal that the uncertainty-based method can provide better fairness performance across datasets, fairness metrics, and foundational models, with a marginal impact on accuracy. Our contribution can be summarized as follows:
While most existing studies focus on fair LLM-based ICL with serialized tabular data, we provide, to our knowledge, the first investigation into preprocessing methods for fair prediction in ICL using transformer-based tabular foundation models.
We propose and investigate three pre-processing intervention methods to enforce fair ICL. These methods aim to reduce the information about the sensitive attributes of in-context samples. We demonstrate that uncertainty-based in-context sample selection can significantly improve the fairness of ICL predictions with a slight drop in accuracy.
We perform extensive experiments on a broad range of start-of-the-art fairness benchmarks and provide insights into contexts where a given fairness intervention performs best in terms of fairness accuracy tradeoff.
We release the code to ease the reproduction of the results and help researchers and practitioners integrate the proposed methods.
We consider a classification task with the giving training data where is an input feature vector, is the corresponding class label, and the corresponding demographic group. The goal is to learn a classifier to accurately predict the target given a sample while being fair w.r.t. demographic information . Several metrics have been proposed to measure fairness at the group or individual levels (Dwork et al., 2012). In this work, we focus on group fairness notions, measuring the performance disparity across different demographic groups, i.e., demographic parity, equalized odds, and equal opportunity (Hardt et al., 2016). A detailed description of these fairness metrics can be found in Appendix B.1.
2 Related works
Numerous methods have been proposed to enforce group fairness in classical machine learning models (Mehrabi et al., 2022; Mbiazi et al., 2023). Model-agnostic methods generally consist of modifying or reweighting the input data (pre-processing) to reduce information about the sensitive attribute (Madras et al., 2018; Creager et al., 2019; Kenfack et al., 2023; Kamiran & Calders, 2012; Celis et al., 2020; Balunović et al., 2021; Feldman et al., 2015; Kenfack et al., 2021) or altering the prediction outcomes (post-processing) to satisfy the desired fairness outcome (Hardt et al., 2016; Petersen et al., 2021). Some other methods adjust the training loss (in-processing) to involve fairness constraints during the training (Agarwal et al., 2018; Zhang et al., 2018; Roh et al., 2020). In contrast to prior works, this work focuses on pre-processing fairness intervention to ensure fair ICL predictions where the downstream model is a foundation model that makes predictions without model updates. We also note that model-agnostic fairness methods, such as pre-processing and post-processing, can be applied to ICL as they do not require model update. Benchmarking their performance on tabular foundation models under the ICL paradigm remains an open research question.
In-context learning with tabular foundation models presents a notable advantage over traditional machine learning approaches by enabling models to adapt dynamically to new data without the need for retraining (Hollmann et al., 2022; Qu et al., 2025; Hollmann et al., 2025). Conventional ML methods typically depend on predefined training datasets, meaning that any alteration in the data or task necessitates a time-consuming and resource-intensive retraining process. In contrast, tabular foundation models utilize in-context learning to execute tasks based on the specific context of the data provided at inference time. This allows these models to interpret and process new tabular data with minimal prior preparation, facilitating more flexible and efficient decision-making (Hollmann et al., 2022). The advantages of this approach are particularly apparent in scenarios where data distributions change over time or when models must quickly adjust to various data tasks without undergoing retraining. Thus, as in-context learning emerges as a powerful tool for real-time, adaptive predictions in complex and dynamic environments, assessing and mitigating biases in the prediction can make its use more socially acceptable.
Studies on fair ICL use large language models with tabular datasets serialized into text or sentences (Bhaila et al., 2024; Hu et al., 2024; Ma et al., 2023). Specifically, Hu et al. (2024) investigates different group sampling methods to improve fairness on ICL predictions and suggests that prioritizing minority group demonstrations can improve the fairness performance of LLMs. Similarly, Bhaila et al. (2024) propose a data augmentation method to mitigate bias and select demonstration examples that can improve the fairness of ICL inference with LLMs. However, LLMs are not specifically trained to handle tabular data, and the overall performance of LLM-based ICL still lags behind well-tuned numeric models such as tree-based models (Hollmann et al., 2025). To our knowledge, this work represents the first initiative that investigates pre-processing fairness intervention on tabular foundation models. Robertson et al. (2024) proposed FairPFN, an in-processing method to improve counterfactual fairness (Kusner et al., 2017) at inference time. The proposed method trains the foundation model with synthetic fair datasets to eliminate the causal dependence on the sensitive attributes. During the pre-training stage, fairness is enforced by predicting the class label of the fair dataset using the biased data as context information. However, this method is computationally expensive since it requires retraining the foundation model. In contrast, we focus on pre-processing methods that do not require model update and aim to achieve group fairness metric instead of counterfactual fairness.
3 Methodology
This section presents three pre-processing techniques proposed in this work to ensure fairer ICL inference on tabular data. In particular, we consider correlation remover (Feldman et al., 2015), group-balanced demonstration selection, and uncertainty-based demonstration selection.
3.1 In-context Samples Transformation
Correlation remover (Feldman et al., 2015; Bird et al., 2020) is a preprocessing method that reduces the correlation between the sensitive and non-sensitive attributes before fitting the model. More specifically, a linear transformation is applied to each non-sensitive feature to reduce its correlation with the sensitive feature. We use the correlation remover as a preprocessing step over the training (in-context example) and testing sets before performing in-context prediction. Ultimately, transforming input features to reduce their linear dependency on the sensitive feature can reduce the reliance on sensitive features in the downstream models. However, as we will see in the results, nonlinear and complex downstream models can still infer the nonlinear dependencies over the sensitive feature and provide unfair results. A more detailed description of correlation remover is provided in Appendix B.2.
3.2 In-context Samples Selection
In this work, we posit that in-context sample selection can have a significant impact on the fairness of ICL prediction. We analyze several demonstration selection methods that can improve the fairness of ICL predictions without any fairness intervention.
Group-balanced demonstration set selection.
Representation bias is a common source of bias in machine learning models (Mehrabi et al., 2022). It occurs when the collected training data does not reflect the demographic diversity of the population. As a result, some demographic subgroups are under-represented, if not represented at all. Recent studies have demonstrated the benefits of group-balanced training data on the fairness properties of the downstream model. Several methods have been proposed to mitigate representation bias in the data, including subsampling the majority group or reweighting the training data based on group proportions (Kamiran & Calders, 2012; Celis et al., 2020). In this paper, we focus on subsampling since current tabular foundation models do not handle sample weights (Hollmann et al., 2025; Qu et al., 2025). Specifically, we perform ICL with a group-balanced demonstration set sampling from each group uniformly at random. When the demonstration set size does allow equal group representation, we subsample the majority group at random. A similar strategy is employed by (Hu et al., 2024) to select demonstrations for few-shot ICL prediction with LLMs. In this paper, we evaluate the effectiveness of this fairness intervention with tabular foundation models instead of using LLMs on serialized tabular data.
Uncertainty-based demonstration set selection
Kenfack et al. (2024b) demonstrated that models trained without fairness constraints can have better fairness properties when the training data consists of samples with uncertain sensitive attributes. We hypothesize that the uncertainty of the sensitive attribute prediction can be a good measure to select demonstrations that improve the fairness in-context predictions. To validate this, we measure the uncertainty of predicting the sensitive attribute in the demonstration set and use samples with high uncertainty for in-context learning. We focus on conformal prediction (Shafer & Vovk, 2008; Vovk et al., 2005) as uncertainty measure since it provides strong theoretical guarantees for the coverage. Instead of returning a single label, a conformal predictor returns a prediction set containing the true label with a probability of at least , with being a user-defined coverage parameter of the conformal prediction (Angelopoulos et al., 2023). For example, setting ensures the prediction set contains the true sensitive attribute value with at least probability. Specifically, samples with prediction sets containing more than one value are uncertain. Intuitively, the coverage parameter controls the fairness-accuracy tradeoff, with meaning no fairness intervention where all the datapoints are used and meaning maximal fairness intervention where only uncertain samples are included in the in-context examples.
Since conformal prediction is model agnostic, we considered both classical methods, e.g., Logistic Regression (LR), and foundation models, e.g., TabPFN for training the sensitive attribute classifier to measure the prediction uncertainties. Note that the method could be applied using other uncertainty measures, such as Monte Carlo dropout and confidence interval (Kenfack et al., 2024b). We focus on conformal prediction due to its rigorous theoretical guarantees, and it does not require a hyperparameter to threshold the level from which a prediction is considered uncertain (Angelopoulos et al., 2023).
3.3 In-context prediction
After performing demonstration selection or transformation using the fairness intervention methods presented previously, we pass them through the tabular foundation model as in-context examples for predicting class labels on the test set. In this paper, we consider TabPFN (Hollmann et al., 2025) and TabICL (Qu et al., 2025) as tabular foundation models. These models use Transformer architectures as their backbones and are trained with extensive synthetic datasets generated from structured causal networks. While TabPFN can handle a maximum of 10k samples with 500 features, TabICL can handle up to 500K samples.
4 Experiments
We describe the experimental setup and perform intensive experiments to answer the following research questions: Does correlation remover and group-balanced demonstration selection effectively reduce information about the sensitive attribute and improve the fairness of ICL? What fairness intervention provides a better fairness-accuracy tradeoff? And what foundation tabular (TabPFN vs TabICL) model can provide a better fairness-accuracy tradeoff?
4.1 Experimental Setup
Datasets
We experiment on tasks from the recently proposed folktables (Ding et al., 2022), which contains data extracted from the American Community Survey (ACS) Public Use Microdata Sample (PUMS) (Ding et al., 2022). More specifically, we experiment with the following ACS PUMS tasks:
ACSIncome: The task involves predicting whether an individual’s income exceeds $50,000. The dataset is filtered to include only individuals over the age of 16 who reported working at least 1 hour per week during the past year and earned a minimum of $100.
ACSMobility: This task involves predicting whether an individual had the same residential address one year ago. The dataset is filtered to include individuals aged between 18 and 35. This filtering increases the difficulty of the task, as more than 90% of the general population tends to stay at the same address year-to-year.
ACSTravelTime: This task predicts whether an individual has a commute longer than 20 minutes. The dataset is filtered to include only employed individuals above the age of 16. The 20-minute threshold corresponds to the median commute time in the US, according to the 2018 ACS PUMS data.
ACSEmployment: The objective is to predict whether an individual is employed, using a dataset filtered to include individuals aged between 16 and 90.
ACSPublicCoverage: The goal is to predict whether an individual has public health insurance. The dataset is filtered to include individuals under 65 years of age and those with an income below $30,000, focusing on low-income individuals who are ineligible for Medicare.
These tasks were selected to reflect a range of real-world predictive challenges with fairness concerns. We use the data of the year 2018 from the state of Alabama (AL), which is one of the states with the largest fairness violation (Ding et al., 2022)222We also perform experiments on data from other states, and observed that the results presented in the paper remain consistent..
A limitation of the ACS PUMS datasets is that they are US-centric; we diversify the experimental setup by including other tasks and datasets. Specifically, we also experiment on the following tabular datasets and tasks:
Diabetes (Gardner et al., 2023): The diabetes prediction task uses features related to physical health, lifestyle factors, and chronic conditions, derived from the BRFSS questionnaires. Demographic attributes like race, sex, state, and income are also included. The target is a binary indicator of whether the respondent has ever been diagnosed with diabetes.
German Credit (Frank, 2010): The German Credit dataset contains 20 attributes of 1,000 individuals. We create the task of classifying people according to whether they have a good or bad credit risk using age (over or below 25 years old) as the sensitive attribute.
CelebA (Liu et al., 2018): The dataset contains 202,599 samples described with 40 facial attributes of human annotated images. We create the task of predicting attractiveness with facial attributes using gender as the sensitive attribute (Kenfack et al., 2024b). Note that we do not train the model with images and consider this task to diversify the experimental tasks.
More details about each dataset, including the sensitive attributes used for fairness evaluation, the number of samples, and the number of features, can be found in the Appendix in Table 1.
Metrics
In addition to the classic accuracy, we consider three group fairness metrics, i.e., Demographic Parity (DP), Equal Opportunity (EOP), and Equalized Odds (EOD), more details can be found in B.1.
Evaluation
For evaluation, we split each dataset into 80%-20%, where the 20% is held out and used to train the sensitive attribute classifier for uncertainty quantification. For the remaining 80% we use 5-fold cross-validation with random shuffle across ten independent runs. This ensures our evaluation is robust and reliable since every data point is used as a test or in-context example across k-fold evaluations. We report the average and standard deviation of fairness and accuracy performance across the k-fold test sets and the ten random independent seeds. As aforementioned, we use TabPFN and TabICL as foundation models for ICL prediction.
Baselines
We measure and compare the fairness and accuracy performance of ICL under four different in-context selection methods:
Vanilla: the vanilla method performs ICL using randomly selected in-context examples from the training data without any fairness consideration.
Balanced: group-balanced in-context examples selection where context examples are randomly drawn with equal demographic group ratio. The majority group is uniformly downsampled across k-fold evaluations and independent runs.
Correlation Remover: in-context transformation with correlation remover (Feldman et al., 2015) where each non-sensitive feature is transformed to reduce correlation with the sensitive ones. We use the fairlearn toolkit (Bird et al., 2020) implementation and fixed the parameter controlling the fairness-accuracy tradeoff to one, meaning maximal fairness.
Uncertain: using the uncertainty of the sensitive attribute prediction to select in-context examples. We use the Mappie implementation of conformal prediction (Cordier et al., 2023) to estimate the uncertainty of the training data. We fix the coverage parameter , of conformal prediction to and only select as in-context example samples with prediction sets equal to two, i.e., samples with uncertain sensitive attributes. We consider two variants of the method under different model classes used to train the sensitive attribute classifier for uncertainty estimation: a variant that uses the traditional Logistic Regression model (Uncertain+LR) and a variant that uses a foundation model (Uncertain+TabPFN).
4.2 Results and discussion
We evaluate several aspects of fairness in ICL prediction with foundational tabular datasets. First, we compare the different baselines considered in terms of fairness and accuracy; For the methods with controllable tradeoff between fairness and accuracy, we vary the hyperparameter controlling the fairness-accuracy tradeoff and compare their Pareto fronts. We then compare the foundation model under the best-performing fairness intervention method. Finally, we provide an ablation study on the impact of the size of the in-context set on fairness and accuracy.
4.2.1 Baseline comparison
Figure 1 summarizes the accuracy and fairness of the different baselines considered across the eight datasets with TabPFN as the foundational model. The group-balanced method shows a marginal improvement in fairness compared to the Vanilla method across datasets. This shows that representation bias is not the only source of bias in the data, and having a group-balanced in-context sample set does not necessarily reduce performance disparity across groups. The unfairness remaining could be due to other sources of bias such as historical bias and measurement bias (Mehrabi et al., 2022).
Surprisingly, Correlation Remover exacerbates unfairness in most datasets. For example, on the ACSIncome dataset, demographic parity increases from to . To further investigate the failure case of Correlation Remover, we measure and compare the reconstruction of sensitive attributes by the foundation model before and after applying the preprocessing intervention. Specifically, we perform ICL using the sensitive attribute as the target variable and use the accuracy of the ICL prediction of the sensitive attribute in the test set as a measure of how the foundation model can reconstruct the sensitive attribute after the fairness intervention. We observe that TabPFN and TabICL can fully reconstruct (up to accuracy) the sensitive attribute based on the feature transformation applied by Correlation Remover (see Table 2 in Appendix). This demonstrates that transformer-based models like TabPFN, pretrained on synthetic data with different sample and feature interactions, can uncover the transformation applied to the non-sensitive features at test-time. By transforming every non-sensitive feature based on the sensitive one, Correlation Remover introduces a sensitive information leakage that the foundational model uncovers and relies upon for predicting the target variable, thereby exacerbating unfairness. Further discussion and results can be found in Appendix D, where we also demonstrate that applying the feature transformation only to the training set and leaving the test set unchanged yields better fairness outcomes compared to the Vanilla approach.
On the other hand, the results also demonstrate that Uncertain methods can significantly improve the fairness of the ICL prediction compared to the different baselines. Table 2 in the Appendix shows that Uncertain methods yield the smallest accuracy of ICL predictions of the sensitive attribute across datasets. This indicates that the selected in-context samples do not encode sufficient information about the sensitive attribute that the foundation model can rely upon for inference, thereby reducing unfairness.
In Appendix C we provide an ablation on the size of in-context examples showing that our findings remain consistent across smaller in-context sizes. More specifically, accuracy significantly increases with the size of the in-context examples, while fairness steadily increases and becomes almost constant when the in-context size exceeds 700 examples.
4.2.2 Fairness Accuracy Tradeoff
In the previous experiment, we fixed the parameters and , controlling the tradeoff between fairness and accuracy for the Correlation Remover and the Uncertain methods, respectively. For a better comparison of the fairness accuracy performances, we consider a range of values between [0, 1] for parameters controlling the tradeoff and plot the Pareto front. Figure 2(a) and 2(b) show the Pareto front on the ACSIncome and ACSMobility datasets, respectively, and compare the Vanilla ICL method, Correlation Remover, and the two variants of the Uncertain method. Results on other datasets can be found in the Appendix (Figure 6) and results using TabICL in Figures 9 and 8. Each scatter point in the figure represents a different value of and for the corresponding ICL method. As can be seen, both variants of the Uncertain consistently have better Pareto dominant points while Correlation Remover can perform worse than the Vanilla ICL in terms of fairness, even with different values of . This further provides evidence of bias amplification of Correlation Remover when used with foundation models. We also see that the Uncertain method can consistently control the fairness accuracy trade-off with smaller values of , enforcing better fairness at the expense of accuracy. The slight decrease in accuracy when the Uncertain method is applied is mainly due to the reduced size of the in-context set, since more and more samples with higher certainty of sensitive attributes are not included in the in-context set.
4.2.3 Foundation Model Comparison: TabICL vs. TabPFN
In the previous experiment, we compared Correlation Remover and Uncertain methods in terms of fairness-accuracy tradeoff by varying their parameter controlling the tradeoff. The results demonstrated that the Uncertain methods often provide better Pareto dominant points compared to the Vanilla ICL and Correlation Remover. And the Uncertain using TabPFN for uncertainty measurement (Uncertain+TabPFN) tends to perform better than using Logistic Regression for uncertainty estimation (Uncertain+LR). In this experiment, we apply the same setup and compare the tradeoffs for the different foundation tabular models under the best-performing fairness intervention, i.e., Uncertain+TabPFN. Figure 3(a) and 3(b) show the fairness accuracy Pareto front on the ACSIncome and ACSMobility datasets, respectively.
We can observe that both foundation models tend to have similar fairness performances while TabPFN often provides better accuracy across datasets. This is consistent with the observation made by the TabICL’s authors when comparing with the TabPFN (Qu et al., 2025). We observe similar trends on other datasets, and the results can be found in Figure 10 of the Appendix.
5 Conclusion and future works
In this study, we proposed and analyzed the effectiveness of three preprocessing methods to enhance the fairness of in-context learning (ICL) predictions. Our empirical results, performed on eight fairness benchmarks, posit the uncertainty-based in-context selection method as a strong baseline for improving the fairness of tabular ICL. The key advantages of this method are threefold: (1) it does not require fine-tuning or retraining the foundation model to enforce the desired fairness metrics; (2) it can consistently improve three widely used group fairness metrics; (3) it offers a parameter to control the fairness-accuracy tradeoff. To our knowledge, this is the first work that explores pre-processing fairness intervention on tabular foundation models. We hope this work will trigger more investigations into fair tabular ICL, since in-context learning as a new learning paradigm will be increasingly adopted into decision-making tools. Interesting future research directions include investigating in-processing and post-processing methods and analyzing the effect of distribution shift between in-context and test examples on fairness and accuracy.
Acknowledgement
The authors thank the Digital Research Alliance of Canada and Denvr for computing resources. SEK is supported by CIFAR and NSERC DG (2021-4086) and UA by NSERC DG (2022-04006).
Ethics Statement
This paper explores ways to reduce unfairness in tabular foundation models, emphasizing fair treatment for various groups. We recognize the significance of fairness in machine learning, especially regarding sensitive attributes like race, gender, and socio-economic status. Our research seeks to uncover and tackle potential biases in these models, thereby enhancing transparency, accountability, and inclusivity. While the proposed method uses a sensitive attributes predictor, which could be unlawful in some countries, we emphasized that predicted sensitive values are not used either for training or measuring unfairness. We use the attribute classifier only to quantify uncertainty, and emphasize that this method should not be used for any purpose other than bias measuring or mitigation.
Appendix
Appendix A Datasets
| Dataset | # Features | # Samples | Sensitive Feature | Prediction Task |
| ACSIncome | 10 | 22,268 | Gender | Income $50,000 |
| ACSEmployment | 16 | 47,777 | Gender | Employment status |
| ACSTravelTime | 16 | 19,492 | Gender | Commute time over 20 minutes |
| ACSMobility | 21 | 8,625 | Gender | Residential mobility |
| ACSPublicCoverage | 19 | 18,525 | Gender | Public health insurance coverage |
| CelebA | 39 | 202,599 | Gender | Attractiveness |
| Diabetes | 183 | 38,575 | Race | Prior diabetes diagnose |
| German | 58 | 990 | Age | Credit risk |
Appendix B Background
B.1 Fairness Metrics
In this work, we focus on group fairness notions that measure the performance disparity across different demographic groups. More specifically, we consider the following three widely used group fairness metrics:
Demographic parity (DP): DP enforces equal positive outcome rate for different groups (Dwork et al., 2012) and is defined as follows:
(1) -
Equalized Odds (EOD): EOdds is satisfied when the model makes correct and incorrect predictions at the same rate for different demographic groups (Hardt et al., 2016). The metric enforces equal true positive and false positive rates across groups and is measured as follows;
(2) -
Equalized Opportunity (EOP): In some settings, one can care more about assessing unfairness when the model makes correct predictions. EOP enforces equal true positive rates across groups, i.e., we only consider in Eq. 2, i.e.,
Empirically, we measure each fairness considered, i.e., Demographic Parity (DP), Equal Opportunity (EOP), and Equalized Odds (EOD) as follows.
Where is the indicator function.
Where and measure the difference between the false positive and the true positive rates across groups, respectively, and are empirically measured as follows.
Where and measure the difference between the false positive and the true positive rates across groups, respectively, and are empirically measured as follows.
Since the disparities can be below on some datasets, we scaled the fairness values reported throughout the paper by 100 to make it easier to read.
B.2 Correlation Remover
The Correlation Remover (Feldman et al., 2015) is a preprocessing technique designed to eliminate linear correlations between sensitive attributes and non-sensitive features in a dataset. This method is particularly useful in mitigating biases that may arise due to such correlations, especially when employing linear models.
Considering a classification task with the given training data where is an input feature vector, is the corresponding class label, and the corresponding demographic group.
To apply Correlation Remover, we assume the training data is formulated as follows:
represents the training data matrix containing sensitive and non-sensitive features.
a matrix of the sensitive features. For simplicity, we assumed in this work , which corresponds to a single binary sensitive attribute.
a matrix of non-senstive features such that
The goal of Correlation Remover is to transform into such that is uncorrelated with , while retaining as much information from the original as possible.
For each non-sensitive feature vector (the -th column of ), the algorithm solves the following least squares problem:
where:
is the mean vector of the sensitive features.
is an -dimensional column vector of ones.
is the weight vector that projects the centered sensitive features onto .
After computing the optimal weight vectors for all }, they are assembled into a weight matrix . The transformed non-sensitive features are then obtained by:
This operation effectively removes the linear correlations between and , resulting in that is uncorrelated with the sensitive features.
Correlation Remover introduces a tunable parameter that controls the extent of correlation removal, i.e., (i) corresponds to full removal of linear correlations, thus best possible fairness; (ii) corresponds no transformation, the original data is used; (iii) corresponds to partial removal, balancing between the original and transformed data, thus controlling the fairness accuracy tradeoff. More specifically, the final transformed dataset is computed as:
Note that is derived using , since is dropped after transformation. The convex combination 10 allows practitioners to adjust the fairness accuracy tradeoff based on specific requirements of their application.
Appendix C Ablation on impact in-context sample size
All the previous experiments used the entire training as in-context examples when possible. Note that the current version of TabPFN cannot handle more than samples (Hollmann et al., 2025). As an ablation study, we vary the in-context sample size ([100, 300, 500, 700, 1500, 2000, 2500, 3000, 4000, 5000]) and measure fairness and accuracy using the same experimental setup as in Section 4.2.1. Figure 9 shows that accuracy significantly increases when more samples are added to the in-context set. Unfainness increases slightly before remaining almost constant when the in-context set size exceeds . We observe that Correlation Remover still has the highest unfairness across in-context size and data. On the other hand, Uncertain+TabPFN method consistently has the lowest fairness violation across in-context sizes, demonstrating its robustness to the size of training data points.
Appendix D On the failure of Correlation Remover
In the main paper, we observed that the transformation 10 on both the training and the testing data can exacerbate unfairness in ICL predictions. We hypothesized that the foundation model inferred the sensitive attribute from the linear transformation applied to each non-sensitive feature, resulting in higher unfairness in the ICL predictions. To validate this hypothesis, we perform ICL prediction of the sensitive attribute after the fairness interventions are applied. As can be seen in Table 2, ICL prediction of the sensitive attribute results in 100% accuracy after Correlation Remover is applied. These results suggest the foundation model still relies on the sensitive attribute after correlation removal is performed.
| Dataset | ICL Method | TabPFN | TabICL | ||
| Accuracy | F1 Score | Accuracy | F1 Score | ||
| ACSIncome | Vanilla | ||||
| Balanced | |||||
| Correlation R. | |||||
| Uncertain+LR | |||||
| Uncertain+TabPFN | |||||
| ACSTravelTime | Vanilla | ||||
| Balanced | |||||
| Correlation R. | |||||
| Uncertain+LR | |||||
| Uncertain+TabPFN | |||||
| ACSPublicCoverage | Vanilla | ||||
| Balanced | |||||
| Correlation R. | |||||
| Uncertain+LR | |||||
| Uncertain+TabPFN | |||||
| ACSEmployment | Vanilla | ||||
| Balanced | |||||
| Correlation R. | |||||
| Uncertain+LR | |||||
| Uncertain+TabPFN | |||||
| ACSMobility | Vanilla | ||||
| Balanced | |||||
| Correlation R. | |||||
| Uncertain+LR | |||||
| Uncertain+TabPFN | |||||
| German Credit | Vanilla | ||||
| Balanced | |||||
| Correlation R. | |||||
| Uncertain+LR | |||||
| Uncertain+TabPFN | |||||
| Diabetes | Vanilla | ||||
| Balanced | |||||
| Correlation R. | |||||
| Uncertain+LR | |||||
| Uncertain+TabPFN | |||||
| CelebA | Vanilla | ||||
| Balanced | |||||
| Correlation R. | |||||
| Uncertain+LR | |||||
| Uncertain+TabPFN | |||||
| Dataset | Correlation R. | TabPFN | TabICL | ||
| Accuracy | F1 Score | Accuracy | F1 Score | ||
| ACSIncome | None | ||||
| Train & Test | |||||
| Train | |||||
| ACSTravelTime | None | ||||
| Train & Test | |||||
| Train | |||||
| ACSPublicCoverage | None | ||||
| Train & Test | |||||
| Train | |||||
| ACSEmployment | None | ||||
| Train & Test | |||||
| Train | |||||
| ACSMobility | None | ||||
| Train & Test | |||||
| Train | |||||
| Diabetes | None | ||||
| Train & Test | |||||
| Train | |||||
| German Credit | None | ||||
| Train & Test | |||||
| Train | |||||
| CelebA | None | ||||
| Train & Test | |||||
| Train | |||||
For further verification, we consider a variant of Correlation Remover where we apply the feature transformation (Eq. 10) only to the training dataset, leaving the test data unchanged. Table 3 shows this variant (Eq. 10) significantly reduces the accuracy of ICL predictions. This demonstrates that the foundation model uses the transformation applied to the testing set as a proxy to fully reconstruct the sensitive attribute. Aïvodji et al. (2021) discuss several scenarios where one can apply a data transformation to reduce correlation with the sensitive attribute. More specifically, they considered scenarios where the transformation is applied (1) both to the training and testing dataset, (2) only to the training set, and (3) only to the test set. We evaluated the fairness accuracy tradeoff of applying feature transformation only to the training set. The results in Figure 5 show that this variant can improve fairness compared to applying the transformation to both the training and testing datasets. This further illustrates the ability of the foundation models to reconstruct the sensitive attributes when the transformation is applied to the test set.
Appendix E Supplementary resutls
| Dataset | ICL Method | Accuracy | DP | EOP | EOD |
| ACSIncome | Vanilla | ||||
| Balanced | |||||
| Correlation R. | |||||
| Uncertain+LR | |||||
| Uncertain+TabPFN | |||||
| ACSEmployment | Vanilla | ||||
| Balanced | |||||
| Correlation R. | |||||
| Uncertain+LR | |||||
| Uncertain+TabPFN | |||||
| ACSPublicCoverage | Vanilla | ||||
| Balanced | |||||
| Correlation R. | |||||
| Uncertain+LR | |||||
| Uncertain+TabPFN | |||||
| ACSTravelTime | Vanilla | ||||
| Balanced | |||||
| Correlation R. | |||||
| Uncertain+LR | |||||
| Uncertain+TabPFN | |||||
| ACSMobility | Vanilla | ||||
| Balanced | |||||
| Correlation R. | |||||
| Uncertain+LR | |||||
| Uncertain+TabPFN | |||||
| Diabetes | Vanilla | ||||
| Balanced | |||||
| Correlation R. | |||||
| Uncertain+LR | |||||
| Uncertain+TabPFN | |||||
| German Credit | Vanilla | ||||
| Balanced | |||||
| Correlation R. | |||||
| Uncertain+LR | |||||
| Uncertain+TabPFN | |||||
| CelebA | Vanilla | ||||
| Balanced | |||||
| Correlation R. | |||||
| Uncertain+LR | |||||
| Uncertain+TabPFN |