### v4

# Towards Fair In-Context Learning with Tabular Foundation Models

Abstract
Transformer-based tabular foundation models have recently demonstrated promising in-context learning (ICL) performance on structured data, emerging as competitive alternatives to gradient-boosted trees. However, the fairness implications of this new paradigm remain largely unexplored. We present the first investigation of fairness in tabular ICL, evaluating three recently proposed foundation models—TabPFNv2, TabICL, and TabDPT—on multiple benchmark datasets. To mitigate biases, we explore three pre-processing fairness-enhancing methods: correlation removal (decorrelating input features from the sensitive attribute), group-balanced sample selection (ensuring equal representation of protected groups in context examples), and uncertainty-based sample selection (prioritizing context examples with high sensitive-attribute prediction uncertainty). Our experiments show that the uncertainty-based strategy consistently improves group fairness metrics (e.g., demographic parity, equalized odds, and equal opportunity) with minimal impact on predictive accuracy. We release our code to facilitate reproducibility (https://github.com/patrikken/Fair-TabICL).
1 Introduction
Tabular data, represented in rows and columns, is a data modality widely used for prediction tasks in domains such as finance and healthcare (asuncion2007uci). Tree-based models such as XGboost (chen2015xgboost) and Gradient-Boosted Trees (ke2017lightgbm) have shown the strongest generalization performance on tabular data. Recently, with the emergence of foundation models, Deep Learning (DL) based models have challenged the dominance of tree-based models (hollmann2025accurate). Foundation models are models pretrained on vast datasets, without a specific task in mind, and they can be adapted across a wide range of downstream tasks. Large language models (LLMs) such as GPT-3 are common examples of foundation models, and they have demonstrated emerging capabilities such as in-context learning (ICL) with few labelled data (brown2020language). In-context learning (ICL) has primarily been applied to natural language tasks using large language models (LLMs). For example, in text classification, labeled examples are formatted as textual demonstrations and provided as context to a language model, enabling it to predict the label of a new instance without any parameter updates or fine-tuning (radford2019language; brown2020language). More recently, efforts have extended ICL to tabular data by serializing table rows into text or sentences (hegselmann2023tabllm). However, since LLMs are not pretrained to model the complex structural relationships inherent in tabular data—such as interactions between rows and columns—their performance on large-scale tabular tasks still lags behind tree-based methods (hegselmann2023tabllm).
Alternatively, recent work has proposed foundation models explicitly tailored for tabular data, achieving competitive performance with tree-based models while reducing the need for extensive model selection and hyperparameter tuning. For instance, TabPFN (hollmann2022tabpfn) is a transformer-based model pretrained on synthetic datasets, and its successor, TabPFNv2 (hollmann2025accurate), extends support to larger datasets with up to 10k samples. Similarly, TabICL (qu2025tabicl), also pretrained on synthetic data, scales to datasets with up to 500k samples. These models leverage synthetic pretraining to encode a wide range of statistical priors, enabling effective target inference from in-context examples.
To better reflect the priors found in real-world datasets, other models incorporate real data during pretraining. Tabular Discriminative Pre-trained Transformer (TabDPT) (ma2024tabdpt) is pretrained directly on real-world datasets, while Real-TabPFN (garg2025real) builds on synthetic pretraining with additional fine-tuning on real data. These approaches generally yield improved performance by aligning the learned representations more closely with real-world data distributions.
Given the strong performance and in-context learning capabilities of tabular foundation models, we will likely see widespread adoption in real-world decision-making tasks. This shift could mark a turning point in how tabular data problems are approached. However, the use of ICL-based models in high-stakes domains—such as healthcare, finance, or criminal justice—raises important ethical concerns. In particular, it is critical to assess their potential to perpetuate or even amplify existing social biases. Traditional machine learning models have already been shown to replicate biases present in the data (mehrabi_survey_2022), and recent studies indicate that LLM-based ICL can also produce biased predictions (hu2024strategic; bhaila2024fair). However, these studies rely on serialized representations of tabular data and therefore inherit the limitations of LLMs in handling tabular structures (ma2024tabdpt).
This paper investigates the fairness of ICL prediction using transformer-based tabular foundation models. First, our study reveals, perhaps unsurprisingly, that while these models focus on improving prediction accuracy, they can also amplify bias. Motivated by recent studies on the sensitivity of ICL performance—in terms of fairness and accuracy— to demonstration selection, we aim to address the following research question: What in-context selection/transformation method can improve the fairness of ICL predictions?
In the fairness literature, several metrics have been proposed to measure fairness at the group or individual levels (dwork_fairness_2012). In this work, we focus on widely used group fairness notions, including demographic parity, equalized odds, and equal opportunity (hardt_equality_2016). This group metrics measure the performance disparity across different demographic groups while we acknowledge that other fairness metrics, beyond group fairness, such as individual or counterfactual fairness, could be used depending on the use case. To achieve these fairness notions, several fairness-enhancing methods have been proposed. They are generally grouped into three categories: pre-processing, in-processing, and post-processing (mehrabi_survey_2022) methods. Projecting these categories into the ICL paradigm, pre-processing methods perform demonstration transformation or selection before predicting in context (hu2024strategic). In-processing methods would fine-tune or retrain the foundation model with fairness constraints (robertson2024fairpfn). Post-processing methods would alter the ICL predictions to improve a given fairness metric (hardt_equality_2016). Pre- and Post-processing methods are more computationally friendly since they do not require model updates. This motivates our choice to focus on the pre-processing techniques and leave post-processing interventions for future exploration. More specifically, we propose and investigate three pre-processing fairness interventions: (i) Correlation Remover (feldman2015certifying), a method that alters each input feature to reduce their correlation with the sensitive attribute; (ii) group-balanced111Underlined represents the method’s name throughout the paper and in the results. in-context selection, ensures that the in-context set is group-balanced; (iii) Uncertainty-based in-context selection, estimates the uncertainty of predicting the sensitive attribute of in-context samples and only selects samples with uncertain predictions. We performed intensive experiments on eight fairness benchmark datasets to investigate the effectiveness of each method in terms of fairness and accuracy. Our results reveal that the uncertainty-based method can provide better fairness performance across datasets, fairness metrics, and foundational models, with marginal impact on accuracy. Our contribution can be summarized as follows:
While most existing studies focus on fair ICL with serialized tabular data, we provide, to our knowledge, the first investigation into preprocessing methods for fair prediction in ICL using transformer-based tabular foundation models.
We propose and investigate three pre-processing intervention methods to enforce fair ICL predictions. These methods aim to reduce the information about the sensitive attributes of in-context samples. We demonstrate that uncertainty-based in-context sample selection can significantly improve the fairness of ICL predictions with a slight drop in utility, e.g., accuracy.
We perform extensive experiments on a broad range of start-of-the-art fairness benchmarks and provide insights into contexts where a given fairness intervention performs best in terms of fairness-utility tradeoff.
We release the code to ease reproduction of the results and help researchers and practitioners integrate the proposed methods.
2 Related works
Fairness.
Numerous methods have been developed to enforce group fairness in classical machine learning models (mehrabi_survey_2022; mbiazi2023survey; kenfack2024a). These methods are often categorized as pre-processing, in-processing, or post-processing approaches. Model-agnostic methods in the pre-processing category typically modify or reweight the input data to reduce information correlated with sensitive attributes (madras_learning_2018; creager_flexibly_2019; kamiran_data_2012; celis2020data; balunovic2021fair; feldman2015certifying). In contrast, post-processing techniques adjust the model’s prediction outcomes after training to satisfy fairness constraints (hardt_equality_2016; petersen2021post). Finally, in-processing approaches embed fairness constraints directly into the training objective (agarwal_reductions_2018; zhang2018mitigating; roh2020fairbatch). Unlike prior work, our approach focuses on pre-processing interventions applied in in-context learning (ICL) settings, where downstream predictions are made by foundation models without any model updates. We emphasize that model-agnostic methods—particularly pre- and post-processing—are especially suitable in ICL because they do not rely on access to or retraining of the model. However, their effectiveness in this setting, especially for tabular foundation models, remains largely unexplored and is the focus of our evaluation.
Tabular Foundation Models.
In-context learning with tabular foundation models presents a notable advantage over traditional machine learning approaches by enabling models to adapt dynamically to new data without the need for retraining (hollmann2022tabpfn; qu2025tabicl; hollmann2025accurate). Conventional ML methods typically depend on predefined training datasets, meaning that any alteration in the data or task necessitates a time-consuming and resource-intensive retraining process. In contrast, tabular foundation models utilize in-context learning to execute tasks based on the specific context of the data provided at inference time. This allows these models to interpret and process new tabular data with minimal prior preparation, facilitating more flexible and efficient decision-making (hollmann2022tabpfn). The advantages of this approach are particularly apparent in scenarios where data distributions change over time or when models must quickly adjust to various data tasks without undergoing retraining. Thus, as in-context learning emerges as a powerful tool for real-time, adaptive predictions in complex and dynamic environments, assessing and mitigating biases in the prediction can make its use more socially acceptable. Existing models are pre-trained using synthetic (qu2025tabicl; hollmann2022tabpfn) or real-world data (ma2024tabdpt). Pre-training on real-world data often provides competitive or better performance, and the performance of synthetic pre-trained tabular foundation models can be boosted with continued pre-training on real-world data (garg2025real). In this work, we consider tabular foundation models pretrained on synthetic (hollmann2025accurate; qu2025tabicl) and real-world data (ma2024tabdpt) and assess the fairness implications of each pretraining strategy.
Fairness in ICL.
Fairness in in-context learning has primarily been studied in the context of large language models (LLMs) applied to tabular data serialized as text(bhaila2024fair; hu2024strategic; ma2023fairness). For instance, hu2024strategic explore group-aware sampling strategies, finding that prioritizing minority group demonstrations improves fairness outcomes. Similarly, bhaila2024fair propose a data augmentation technique that guides demonstration selection to reduce bias during inference. While related in spirit, our approach differs in two key ways: (1) we focus on numerical tabular foundation models (rather than LLMs), and (2) our uncertainty-based selection method aims to reduce model reliance on sensitive attributes rather than optimize informativeness per se. Prior uncertainty-driven methods(mavromatis2023examples; kung2023active) focus on selecting informative examples under a labeling budget, whereas we use uncertainty to guide fair demonstration selection. It is also important to highlight that LLMs, though flexible, are not optimized for tabular data and often perform worse than specialized numeric models, such as gradient-boosted trees (hegselmann2023tabllm). As such, our work fills a key gap by investigating fairness interventions tailored to numeric tabular foundation models. To our knowledge, this is the first study to evaluate pre-processing fairness methods in this emerging model class.
Fairness-Aware Tabular Foundation Models.
Recent work by robertson2024fairpfn introduced FairPFN, a TabPFN-like model trained to suppress the causal effect of sensitive attributes during pretraining. Their approach seeks counterfactual fairness (kusner2017counterfactual), ensuring that model predictions remain invariant when sensitive attributes are counterfactually changed. Our work is conceptually distinct in two respects. First, we do not require model retraining and rely on model-agnostic pre-processing methods, making our approach broadly applicable to any pretrained tabular foundation model. Second, we aim for group fairness, focusing on improving performance disparities between subgroups, rather than enforcing counterfactual invariance at the individual level.
3 Methodology
Problem Setup
We consider a classification task with the given training data where is an input feature vector, is the corresponding class label, and the corresponding demographic group. The goal is to obtain a classifier , via ICL, to accurately predict the target given a sample while being fair w.r.t. demographic information . Several metrics have been proposed to measure fairness at the group or individual levels (dwork_fairness_2012). In this work, we focus on group fairness notions, measuring the performance disparity across different demographic groups, i.e., demographic parity, equalized odds, and equal opportunity (hardt_equality_2016). A detailed description of these fairness metrics can be found in Appendix B.1.
This section presents three pre-processing techniques proposed in this work to ensure fairer ICL inference on tabular data. In particular, we consider correlation remover (feldman2015certifying), group-balanced demonstration selection, and uncertainty-based demonstration selection.
3.1 In-context Samples Transformation
Correlation remover (feldman2015certifying; bird2020fairlearn) is a preprocessing method that reduces the correlation between the sensitive and non-sensitive attributes before fitting the model. More specifically, a linear transformation is applied to each non-sensitive feature to reduce its correlation with the sensitive feature. We use the correlation remover as a preprocessing step over the training (in-context example) and testing sets before performing in-context prediction. Ultimately, transforming input features to reduce their linear dependency on the sensitive feature can reduce the reliance on sensitive features in the downstream models. However, as we will see in Section 4.5, nonlinear and complex downstream models can still infer the nonlinear dependencies over the sensitive feature and provide unfair results. A more detailed description of correlation remover can be found in Appendix B.2.
3.2 In-context Samples Selection
In this work, we posit that in-context sample selection can have a significant impact on the fairness of ICL prediction. We analyze two demonstration selection methods that can improve the fairness of ICL predictions without model update.
3.2.1 Group-balanced demonstration set selection.
Representation bias is a common source of bias in machine learning models (mehrabi_survey_2022). It occurs when the collected training data does not reflect the demographic diversity of the population. As a result, some demographic subgroups are under-represented, if not represented at all. Recent studies have demonstrated the benefits of group-balanced training data on the fairness properties of the downstream model. Several methods have been proposed to mitigate representation bias in the data, including subsampling the majority group or reweighting the training data based on group proportions (kamiran_data_2012; celis2020data). In this paper, we focus on subsampling since current tabular foundation models do not handle sample weights (hollmann2025accurate; qu2025tabicl). We perform ICL with a group-balanced demonstration set sampling from each group uniformly at random. When the demonstration set size does allow equal group representation, we subsample the majority group at random. A similar strategy is employed by (hu2024strategic) to select demonstrations for few-shot ICL prediction with LLMs. In this paper, we evaluate the effectiveness of this fairness intervention with tabular foundation models instead of using LLMs on serialized tabular data.
3.2.2 Uncertainty-based demonstration set selection.
kenfack2024fairness demonstrated that models trained without fairness constraints can have better fairness properties when the training data consists of samples with uncertain sensitive attributes. We hypothesize that the uncertainty of the sensitive attribute prediction can be a good measure to select demonstrations that improve the fairness of in-context predictions. To validate this, we measure the uncertainty of predicting the sensitive attribute in the demonstration set and use samples with high uncertainty for in-context learning. We focus on conformal prediction (shafer2008tutorial; vovk2005algorithmic) as uncertainty measure since it provides strong theoretical guarantees for the coverage. Instead of returning a single label, a conformal predictor returns a prediction set containing the true label with a probability of at least , with being a user-defined coverage parameter of the conformal prediction (angelopoulos2023conformal). For example, setting ensures the prediction set contains the true sensitive attribute value with at least probability. Specifically, samples with prediction sets containing more than one value are uncertain. Intuitively, the coverage parameter controls the fairness-utility tradeoff, with meaning no fairness intervention where all the datapoints are used and meaning maximal fairness intervention where only uncertain samples are included in the in-context examples. We show in Section 4.2.1 how the coverage parameter of the conformal predictor consistently controls the tradeoff between accuracy and fairness across fairness metrics and downstream foundational models.
Since conformal prediction is model agnostic, we considered both classical methods, e.g., Logistic Regression (LR), and foundation models, e.g., TabPFN for training the sensitive attribute classifier to measure the prediction uncertainties. Note that the method could be applied using other uncertainty measures, such as Monte Carlo dropout and confidence interval (kenfack2024fairness). We focus on conformal prediction due to its rigorous theoretical guarantees, and it does not require a hyperparameter to threshold the level from which a prediction is considered uncertain (angelopoulos2023conformal). More details about uncertainty measurement with conformal prediction can be found in Appendix B.3.
3.3 In-Context Prediction
After performing demonstration selection or transformation using the fairness intervention methods presented previously, we pass them through the tabular foundation model as in-context examples for predicting class labels on the test set. In this paper, we consider three tabular foundation models: TabDPT (ma2024tabdpt), TabICL (qu2025tabicl), and TabPFN v2 (hollmann2025accurate). While all these models use Transformer architectures as their backbones and are pretrained on a large amount of datasets, TabICL and TabPFN use only synthetic data generated from structured causal networks, and TabDPT uses real-world data. The current version of TabPFN can handle a maximum of 10k samples with 500 features, and TabICL can handle up to 500K samples. We randomly subsample in-context examples when the context set size exceeds the maximum number of samples allowed by the foundation model being evaluated.
4 Experiments
We describe the experimental setup and perform intensive experiments to answer the following research questions:
(R1) Does group-balanced demonstration selection and correlation remover effectively reduce information about the sensitive attribute and improve the fairness of ICL?
(R2) What fairness intervention provides a better fairness-utility tradeoff?
(R3) What foundation tabular model can provide a better fairness-utility tradeoff?
4.1 Experimental Setup
Datasets
We experiment on tasks from folktables (ding_retiring_2022), which contains data extracted from the American Community Survey (ACS) Public Use Microdata Sample (PUMS). More specifically, we experiment with the following ACS PUMS tasks: ACSIncome, ACSMobility, ACSTravelTime, ACSEmployment, ACSPublicCoverage. These tasks reflect a range of real-world predictive challenges with fairness concerns. We use the data of the year 2018 from the state of Alabama (AL), which is one of the states with the largest fairness violation (ding_retiring_2022). A limitation of the ACS PUMS datasets is that they are US-centric; we diversify the experimental setup by including other tasks and datasets. Specifically, we also experiment on the other tabular datasets and tasks including: Diabetes (gardner2023tableshift), German Credit (frank2010uci), and CelebA (liu2018large). More details about each dataset, including the sensitive attributes used, the number of samples, and the number of features, can be found in the Appendix A.
Metrics
We measure the utility of the model using accuracy, and measure fairness using three group fairness metrics, i.e., Demographic Parity (DP), Equal Opportunity (EOP), and Equalized Odds (EOD). More details about group fairness metrics can be found in B.1.
Baselines
We evaluate ICL under four in-context selection strategies, comparing both fairness and accuracy:
Vanilla: randomly selects in-context examples from the training set without fairness considerations.
Group Balanced: samples in-context examples to maintain equal group ratios, uniformly downsampling the majority group across runs.
Correlation Remover (CR): applies correlation-removal (feldman2015certifying) to reduce dependence between sensitive and non-sensitive features in both in-context and test examples. We use the fairlearn implementation (bird2020fairlearn) with varying values of , where enforces maximal fairness.
Uncertain: selects in-context examples with high uncertainty in sensitive attribute prediction. We estimate uncertainty using conformal prediction (Cordier_Flexible_and_Systematic_2023) as implemented in Mapie (taquet2022mapie), varying the coverage parameter to control the fairness–utility tradeoff. We study two variants of sensitive attribute classifier: one with a Logistic Regression classifier (Uncertain+LR) and one with a foundation model (Uncertain+TabPFN).
Evaluation
For evaluation, we held out 20% of the data to train the sensitive attribute classifier used in the uncertainty-based baseline. The remaining data were split into 80% for training and 20% for testing. We applied the fairness interventions to the training set to derive in-context samples, and reported fairness and accuracy on the test set, averaged over three random seeds. This setup provides robustness, as each data point can appear as a test sample or an in-context sample across seeds. As noted above, we use TabICL, TabDPT, and TabPFN with their default parameters.
4.2 Results and Discussion
We evaluate fairness in ICL predictions with tabular foundation models along several dimensions. First, we compare the baseline methods in terms of fairness and accuracy. For interventions with a tunable fairness–utility tradeoff, we vary the corresponding hyperparameter and compare their Pareto fronts. While we use accuracy as the main utility metric, we verify in the Appendix that results are consistent when using ROC AUC. Next, we compare foundation models under the best-performing intervention. Finally, we conduct an ablation on the effect of in-context set size and analyze the failure case of the correlation-removal method.
4.2.1 Fairness-Utility Tradeoff
Figure 1 shows the fairness–utility Pareto fronts on ACSIncome, ACSMobility, and ACSTravelTime datasets using TabPFN as the foundation model. We compare the Vanilla baseline, Group Balanced, Correlation Remover, and the two variants of our uncertainty-based method (Uncertain+LR and Uncertain+TabPFN). For Correlation Remover and Uncertain, we vary the tradeoff parameters and over , with each shaded point corresponding to a parameter setting. The main findings, consistent across datasets and models, are as follows:
(R1) The Group Balanced method yields only marginal fairness gains over Vanilla ICL, with similar utility. This suggests that representation bias alone does not explain observed disparities; even with group-balanced contexts, subgroup performance gaps remain, likely due to other sources of bias such as historical, algorithmic, or measurement bias (mehrabi_survey_2022).
(R2) In contrast, both variants of Uncertain consistently achieve Pareto-dominant points. They offer stable control over the fairness–accuracy tradeoff across values, enabling significant fairness improvements with modest accuracy loss. The reduced accuracy stems mainly from smaller in-context sets, since excluding low-uncertainty samples decreases the size of the in-context set, which can reduce accuracy as we discuss in Section 4.3).
(R2, Uncertainty variants) Among the two uncertainty-based methods with different model classes, Uncertain+TabPFN typically yields stronger Pareto fronts, indicating that TabPFN provides more reliable conformal predictions than logistic regression, leading to better fairness-aware sample selection.
4.2.2 Accuracy of sensitive attribute reconstruction
| Dataset | ICL Method | TabPFN | TabICL | ||
| Accuracy | F1 Score | Accuracy | F1 Score | ||
| ACSIncome | Vanilla | ||||
| Group Balanced | |||||
| Correlation R. | |||||
| Uncertain+LR | |||||
| Uncertain+TabPFN | |||||
| ACSTravelTime | Vanilla | ||||
| Group Balanced | |||||
| Correlation R. | |||||
| Uncertain+LR | |||||
| Uncertain+TabPFN | |||||
| ACSPublicCoverage | Vanilla | ||||
| Group Balanced | |||||
| Correlation R. | |||||
| Uncertain+LR | |||||
| Uncertain+TabPFN | |||||
| ACSEmployment | Vanilla | ||||
| Group Balanced | |||||
| Correlation R. | |||||
| Uncertain+LR | |||||
| Uncertain+TabPFN | |||||
| ACSMobility | Vanilla | ||||
| Group Balanced | |||||
| Correlation R. | |||||
| Uncertain+LR | |||||
| Uncertain+TabPFN | |||||
| German Credit | Vanilla | ||||
| Group Balanced | |||||
| Correlation R. | |||||
| Uncertain+LR | |||||
| Uncertain+TabPFN | |||||
| Diabetes | Vanilla | ||||
| Group Balanced | |||||
| Correlation R. | |||||
| Uncertain+LR | |||||
| Uncertain+TabPFN | |||||
| CelebA | Vanilla | ||||
| Group Balanced | |||||
| Correlation R. | |||||
| Uncertain+LR | |||||
| Uncertain+TabPFN | |||||
We assess whether foundation models can reconstruct the sensitive attribute after different fairness interventions. Specifically, we perform ICL to predict the sensitive attribute and measure test accuracy. Effective debiasing should lower this accuracy, ideally approaching random guessing ().
Table 1 shows that the Group Balanced intervention only marginally reduces reconstruction accuracy for TabPFN and TabICL on ACS datasets. This limited effect is expected, since ACS is nearly group-balanced for gender (ding_retiring_2022), e.g., for the state of Alabama, the gender distribution is females, males. In contrast, the Diabetes dataset is highly imbalanced by race, and group balancing reduces reconstruction accuracy from for the vanilla baseline to .
For Correlation Remover, we observe the opposite effect: reconstruction accuracy rises to nearly . We identify this as the source of bias amplification. Transformer-based foundation models, pretrained on synthetic data with diverse structural priors, appear to detect and exploit the transformations applied by Correlation Remover. Because each non-sensitive feature is modified using the sensitive attribute, sensitive information leaks into the transformed features. The models leverage this leakage, relying on hidden sensitive signals to predict the target variable and thereby amplifying unfairness. We provide further analysis in Section 4.5, showing that applying correlation removal only to the training set—while keeping the test set unchanged—improves fairness compared to the Vanilla baseline.
Finally, the Uncertain methods yield the lowest reconstruction accuracy across datasets. This suggests that the selected in-context samples carry little information about the sensitive attribute, limiting the model’s ability to exploit it and thereby improving fairness.
4.2.3 Comparison of Foundation Models: TabICL vs. TabDPT vs. TabPFN
In the previous experiment, we compared the Correlation Remover and Uncertain methods in terms of their fairness–accuracy tradeoffs by varying their respective control parameters. The results showed that the Uncertain methods often achieved better Pareto-dominant points compared to both Vanilla ICL and Correlation Remover. Notably, using TabPFN for uncertainty estimation (Uncertain+TabPFN) outperformed the variant that relies on logistic regression (Uncertain+LR).
In this experiment, we extend the comparison to different foundation models by applying the same setup using our best-performing fairness intervention: Uncertain+TabPFN. Figures 2(a) and 2(b) show the fairness–accuracy Pareto fronts on the ACSIncome and ACSMobility datasets, respectively.
Without any fairness intervention (i.e., using the vanilla ICL approach), there is no clear winner among the foundation models in terms of fairness; the results vary across datasets. While one might expect TabDPT —pretrained on real-world data—to encode more real-world biases and thus exhibit poorer fairness performance, this is not consistently observed. In fact, TabDPT often performs comparably to or better than foundation models pretrained on synthetic data (R3).
Under fairness intervention with Uncertain+TabPFN, all three foundation models exhibit similar fairness performance, demonstrating that the Uncertain method can consistently control the fairness–accuracy trade-off regardless of the underlying foundation model. However, both TabPFN and TabDPT consistently achieve higher accuracy across datasets (R3). These findings are consistent with previous studies: the benchmark erickson_tabarena_2025 reported strong performance for TabPFN (qu2025tabicl), while the authors of TabDPT highlighted its competitive performance relative to TabPFN. We observe similar patterns across additional datasets, and results are provided in the Appendix (see Figure 5).
4.3 Ablation on Impact In-context Sample Size
In all previous experiments, we used the full training set as the in-context example set whenever possible. For example, the current version of TabPFN is limited to handling a maximum of 10,000 samples (hollmann2025accurate). To understand the impact of in-context set size, we conduct an ablation study by varying the number of in-context examples across the range [100,300,500,700,1500,2000,2500,3000,4000,5000], while keeping the evaluation setup identical as described above.
Figure 3 shows that accuracy increases substantially with larger in-context sets, particularly in the lower range. However, unfairness shows a slight increase initially and then stabilizes once the in-context set exceeds approximately 700 examples. Across all in-context sizes and datasets, the Correlation Remover method consistently exhibits the highest level of unfairness, reinforcing earlier observations of its bias amplification. In contrast, the Uncertain+TabPFN method consistently achieves the lowest fairness violation, demonstrating robustness to the size of the in-context set. This indicates that even when only a subset of training data can be used, Uncertain+TabPFN remains effective at mitigating bias without overly compromising accuracy. We observe a similar trend when using TabICL model, and results are provided in Appendix 13.
In Appendix C, we perform a computational analysis comparing running and inference time of different fainress interventions and foundation models. This analysis shows that fairness interventions themselves are computationally inexpensive. The runtime is dominated by the choice of the tabular foundation model. TabPFN is the most efficient, benefiting from its compact architecture. The uncertainty-based variant with LR is fastest, while using it with TabPFN itself for uncertainty estimation adds moderate overhead..
4.4 Comparison with LLM-based ICL Methods
| Method | Foundation model | Accuracy | ||
| Vanilla | TabPFN | |||
| Vanilla | LLaMA-2-13B | |||
| hu2024strategic | GPT-3.5-turbo | |||
| bhaila2024fair | LLaMA-2-13B | |||
| Uncertain+TabPFN | TabPFN |
Despite LLMs’ in-context learning capabilities, they often lag behind classical machine learning models on large-scale tabular datasets (hegselmann2023tabllm). To validate the effectiveness of our uncertainty-based demonstration selection method, we compare it against fairness-aware LLM-based fair ICL approaches proposed by hu2024strategic and bhaila2024fair on the widely studied Adult dataset (asuncion2007uci). Table 2 presents the results. Vanilla TabPFN achieves an accuracy of , significantly outperforming the LLaMA-2-13B baseline, which reaches only . This 10-point gap underscores the performance advantage of specialized models for tabular data, consistent with prior findings that LLMs struggle on such inputs when used naively (hegselmann2023tabllm).
While these comparisons rely on LLMs that are not the most recent generation, and performance may have improved with newer models, the results nonetheless highlight the continued strength of tabular foundation models like TabPFN in this setting. Moreover, when fairness interventions are applied, our uncertainty-based selection strategy consistently preserves high predictive performance while achieving comparable or better fairness outcomes than the LLM-based fair ICL approaches—demonstrating that principled sample selection can improve fairness without sacrificing accuracy.
4.5 On the Failure of Correlation Remover
| Dataset | Fairness Intervention | TabPFN | TabICL | ||
| Accuracy | F1 Score | Accuracy | F1 Score | ||
| ACSIncome | None | ||||
| Correlation R. (S1) | |||||
| Correlation R. (S2) | |||||
| ACSTravelTime | None | ||||
| Correlation R. (S1) | |||||
| Correlation R. (S2) | |||||
| ACSPublicCoverage | None | ||||
| Correlation R. (S1) | |||||
| Correlation R. (S2) | |||||
| ACSEmployment | None | ||||
| Correlation R. (S1) | |||||
| Correlation R. (S2) | |||||
| ACSMobility | None | ||||
| Correlation R. (S1) | |||||
| Correlation R. (S2) | |||||
| Diabetes | None | ||||
| Correlation R. (S1) | |||||
| Correlation R. (S2) | |||||
| German Credit | None | ||||
| Correlation R. (S1) | |||||
| Correlation R. (S2) | |||||
| CelebA | None | ||||
| Correlation R. (S1) | |||||
| Correlation R. (S2) | |||||
In our previous experiments, we observed that applying the CR to both the training and test data can exacerbate unfairness in ICL predictions. We hypothesized that the foundation model infers the sensitive attribute from the linear transformation applied to each non-sensitive feature, which inadvertently leaks sensitive information and increases unfairness. To validate this hypothesis, we conducted correlation removal prediction of the sensitive attribute after applying fairness interventions. As shown in Table 1, the ICL prediction of the sensitive attribute achieves accuracy following the application of CR. This indicates that the foundation model continues to rely heavily on the sensitive attribute even after correlation removal is performed.
For further verification, we tested a variant of correlation removal where the feature transformation (see Eq. 10 in Appendix) is applied only to the training data, leaving the test data unchanged (referred to as variant S2). Table 3 shows that this variant significantly reduces the ICL prediction accuracy of the sensitive attribute. This demonstrates that the foundation model exploits the transformation applied to the test set in the original correlation removal method (variant S1) as a proxy to fully reconstruct sensitive attributes.
This observation aligns with prior discussions by aivodji2021local regarding scenarios where data transformations reduce correlation with sensitive attributes. Besides the classical correlation removal setting (S1), where transformations are applied to both training and test sets, we evaluated the fairness–accuracy trade-off when applying the transformation only to the training set (S2).
As illustrated in Figure 4, variant S2 of CR substantially improves fairness compared to S1. This confirms the foundation models’ capacity to reconstruct sensitive information when the test set is transformed, leading to bias amplification. And the drop in reconstruction accuracy of the sensitive attribute directly supports the observed improvement in fairness when using variant S2
Based on these findings, we recommend practitioners apply the S2 variant of Correlation Remover—transforming only the training data—when using correlation removal in in-context learning frameworks, to avoid sensitive attribute leakage and mitigate bias amplification during testing.
5 Conclusion and future works
In this study, we proposed and analyzed the effectiveness of three preprocessing methods to enhance the fairness of in-context learning (ICL) predictions. Our empirical results, performed on eight fairness benchmarks, posit the uncertainty-based in-context selection method as a strong baseline for improving the fairness of tabular ICL. The key advantages of this method are threefold: (1) it does not require fine-tuning or retraining the foundation model to enforce the desired fairness metrics; (2) it can consistently improve three widely used group fairness metrics; (3) it offers a parameter to control the fairness-utility tradeoff. To our knowledge, this is the first work that explores pre-processing fairness intervention on tabular foundation models. We hope this work will trigger more investigations into fair tabular ICL, since in-context learning as a new learning paradigm will be increasingly adopted into decision-making tools. Interesting future research directions include investigating in-processing and post-processing methods and analyzing the effect of distribution shift between in-context and test examples on fairness and accuracy.
Ethics Statement
This paper explores ways to reduce unfairness in tabular foundation models, emphasizing fair treatment for various groups. We recognize the significance of fairness in machine learning, especially regarding sensitive attributes like race, gender, and socio-economic status. Our research seeks to uncover and tackle potential biases in these models, thereby enhancing transparency, accountability, and inclusivity. While the proposed method uses a sensitive attributes predictor, which could be unlawful in some countries, we emphasized that predicted sensitive values are not used either for training or measuring unfairness. We use the attribute classifier only to quantify uncertainty, and emphasize that this method should not be used for any purpose other than bias measuring or mitigation.
Appendix
Appendix A Datasets
We experiment on tasks from the recently proposed folktables (ding_retiring_2022), which contains data extracted from the American Community Survey (ACS) Public Use Microdata Sample (PUMS) (ding_retiring_2022). More specifically, we experiment with the following ACS PUMS tasks:
ACSIncome: The task involves predicting whether an individual’s income exceeds $50,000. The dataset is filtered to include only individuals over the age of 16 who reported working at least 1 hour per week during the past year and earned a minimum of $100.
ACSMobility: This task involves predicting whether an individual had the same residential address one year ago. The dataset is filtered to include individuals aged between 18 and 35. This filtering increases the difficulty of the task, as more than 90% of the general population tends to stay at the same address year-to-year.
ACSTravelTime: This task predicts whether an individual has a commute longer than 20 minutes. The dataset is filtered to include only employed individuals above the age of 16. The 20-minute threshold corresponds to the median commute time in the US, according to the 2018 ACS PUMS data.
ACSEmployment: The objective is to predict whether an individual is employed, using a dataset filtered to include individuals aged between 16 and 90.
ACSPublicCoverage: The goal is to predict whether an individual has public health insurance. The dataset is filtered to include individuals under 65 years of age and those with an income below $30,000, focusing on low-income individuals who are ineligible for Medicare.
These tasks were selected to reflect a range of real-world predictive challenges with fairness concerns. A limitation of the ACS PUMS datasets is that they are US-centric; we diversify the experimental setup by including other tasks and datasets. Specifically, we also evaluate on the following tabular datasets and tasks:
Diabetes (gardner2023tableshift): The diabetes prediction task uses features related to physical health, lifestyle factors, and chronic conditions, derived from the BRFSS questionnaires. Demographic attributes like race, sex, state, and income are also included. The target is a binary indicator of whether the respondent has ever been diagnosed with diabetes.
German Credit (frank2010uci): The German Credit dataset contains 20 attributes of 1,000 individuals. We create the task of classifying people according to whether they have a good or bad credit risk using age (over or below 25 years old) as the sensitive attribute.
CelebA (liu2018large): The dataset contains 202,599 samples described with 40 facial attributes of human annotated images. We create the task of predicting attractiveness with facial attributes using gender as the sensitive attribute (kenfack2024fairness). Note that we do not train the model with images and consider this task to diversify the experimental tasks.
| Dataset | # Features | # Samples | Sensitive Feature | Prediction Task |
| ACSIncome | 10 | 22,268 | Gender | Income $50,000 |
| ACSEmployment | 16 | 47,777 | Gender | Employment status |
| ACSTravelTime | 16 | 19,492 | Gender | Commute time over 20 minutes |
| ACSMobility | 21 | 8,625 | Gender | Residential mobility |
| ACSPublicCoverage | 19 | 18,525 | Gender | Public health insurance coverage |
| CelebA | 39 | 202,599 | Gender | Attractiveness |
| Diabetes | 183 | 38,575 | Race | Prior diabetes diagnose |
| German | 58 | 990 | Age | Credit risk |
Appendix B Background
B.1 Fairness Metrics
In this work, we focus on group fairness notions that measure the performance disparity across different demographic groups. More specifically, we consider the following three widely used group fairness metrics:
Demographic parity (DP): DP enforces equal positive outcome rate for different groups (dwork_fairness_2012) and is defined as follows:
(1) -
Equalized Odds (EOD): EOdds is satisfied when the model makes correct and incorrect predictions at the same rate for different demographic groups (hardt_equality_2016). The metric enforces equal true positive and false positive rates across groups and is measured as follows;
(2) -
Equalized Opportunity (EOP): In some settings, one can care more about assessing unfairness when the model makes correct predictions. EOP enforces equal true positive rates across groups, i.e., we only consider in Eq. 2, i.e.,
Empirically, we measure each fairness considered, i.e., Demographic Parity (DP), Equal Opportunity (EOP), and Equalized Odds (EOD) as follows.
Where is the indicator function.
Where and measure the difference between the false positive and the true positive rates across groups, respectively, and are empirically measured as follows.
Where and measure the difference between the false positive and the true positive rates across groups, respectively, and are empirically measured as follows.
Since the disparities can be below on some datasets, we scaled the fairness values reported throughout the paper by 100 to make it easier to read.
B.2 Correlation Remover
The Correlation Remover (feldman2015certifying) is a preprocessing technique designed to eliminate linear correlations between sensitive attributes and non-sensitive features in a dataset. This method is particularly useful in mitigating biases that may arise due to such correlations, especially when employing linear models.
Considering a classification task with the given training data where is an input feature vector, is the corresponding class label, and the corresponding demographic group.
To apply Correlation Remover, we assume the training data is formulated as follows:
represents the training data matrix containing sensitive and non-sensitive features.
a matrix of the sensitive features. For simplicity, we assumed in this work , which corresponds to a single binary sensitive attribute.
a matrix of non-senstive features such that
The goal of Correlation Remover is to transform into such that is uncorrelated with , while retaining as much information from the original as possible.
For each non-sensitive feature vector (the -th column of ), the algorithm solves the following least squares problem:
where:
is the mean vector of the sensitive features, i.e., is the mean of -th the sensitive feature.
is an -dimensional column vector of ones.
is the weight vector that projects the centered sensitive features onto .
After computing the optimal weight vectors for all }, they are assembled into a weight matrix . The transformed non-sensitive features are then obtained by:
This operation effectively removes the linear correlations between and , resulting in that is uncorrelated with the sensitive features.
Correlation Remover introduces a tunable parameter that controls the extent of correlation removal, i.e., (i) corresponds to full removal of linear correlations, thus best possible fairness; (ii) corresponds no transformation, the original data is used; (iii) corresponds to partial removal, balancing between the original and transformed data, thus controlling the fairness accuracy tradeoff. More specifically, the final transformed dataset is computed as:
Note that is derived using , since is dropped after transformation. The convex combination 10 allows practitioners to adjust the fairness accuracy tradeoff based on specific requirements of their application.
B.3 Uncertainty measure with conformal prediction
While any uncertainty measurement method could be used in our method, we employ conformal prediction due to its strong coverage guarantees. Specifically, we used Split Conformal Prediction (angelopoulos2023conformal), which is a distribution-free and model-agnostic method that provides prediction sets for classification tasks, ensuring a user-specified coverage level with no assumptions beyond data exchangeability.
Given labeled data with binary sensitive attribute we split the data into proper training set ( ) and calibration set (). In the experiments, we did a 50%-50% split of the dataset with sensitive attribute to obtain and .
We then train a probabilistic classifier , on , yielding predictions:
In our setup, could be Logistic Regression (Uncertain+LR) or TabPFN (Uncertain+TabPFN). The starting point for conformal prediction is what is called a nonconformity measure, a real-valued function that measures how a prediction is different from any possible class label.
Nonconformity Scores
After training on the proper training set, we use the calibration dataset to compute the nonconformity scores, which measure how far the prediction is from the true label. More specifically, for each calibration point , we considered the nonconformity score is defined as:
We then compute the quantile threshold based on the user-defined target coverage , e.g., mean 95% coverage. Specifically, is defined based on quantile of the nonconformity scores.
Prediction Set
The quantile threshold is used to build the prediction set of data points from the test set. More specifically, for a test sample , we compute the prediction probability and derive its prediction set as follows:
When the prediction set only contains {0} or {1}, then the prediction is confident with at least probability, while when the prediction set contains both labels, i.e., {0, 1}, the prediction is considered uncertain. Our uncertainty-based demonstration selection method uses only samples whose prediction set contains two values. The coverage guarantee of conformal prediction holds under the assumption that the calibration and test data are exchangeable. We use the open-source implementation of split conformal prediction provided by the MAPIE Python package.
Example.
Consider a simplified task consisting of two non-sensitive features ( and ), one binary target (), and a binary sensitive attribute ().
To estimate the uncertainties, we first train a sensitive attribute classifier, using a fraction of the dataset (20% in our experiments). We train the sensitive attribute classifier (Logistic Regression or TabPFN) using and to predict . A conformal predictor returns for a given (test) sample a prediction set that contains the true sensitive attribute with probability of at least . For example, setting means 95% of the data will contain the true label in their prediction sets. The size of the prediction set, therefore, provides information about the prediction uncertainty, i.e., means the prediction is confident with at least probability, and means the prediction is uncertain. The intuition of our method is that using samples with uncertain predictions makes it challenging for the foundation model to infer and rely on the sensitive attributes to make predictions on the target.
We therefore include a training example in the context set if and filter it out otherwise. helps to control the fairness-utility tradeoff since a smaller corresponds to filtering out more demonstration examples with higher confidence sensitive attributes prediction, which can better improve fairness while impacting accuracy as the size of the in-context set is reduced.
Appendix C Runtime Comparison of Fairness Interventions
In addition to fairness–utility tradeoffs, we provide a detailed comparison of the computational cost of the fairness interventions across tabular foundation models. While prior sections focused on predictive performance, practical deployment also requires understanding efficiency, particularly in large-scale or resource-constrained settings.
Setup.
We measured wall-clock running time on an NVIDIA A100 GPU (20 GB memory). Each run includes preprocessing, fairness intervention, and in-context learning inference using the full training set and test set. Results are averaged over three seeds, with standard deviations reported. The largest dataset in our experiments, CelebA, was used to provide representative estimates of runtime under high load. Table 5 summarizes runtimes across methods and models. We make the following observations: (i) TabPFN is the most efficient (8–47s), benefiting from its compact architecture. The uncertainty-based variant with LR is fastest, while using TabPFN itself for uncertainty estimation adds moderate overhead. (ii) TabICL shows moderate cost (57–97s), with relatively small variance across interventions. (iii) TabDPT is substantially slower ( 595s), reflecting the higher computational demands of its transformer-based architecture.
Across models, uncertainty-based with Logistic Regression is consistently the fastest intervention, while correlation removal and group balancing introduce negligible overhead compared to vanilla inference. Recall TabPFN uses at most 10K samples, which reduces the context size compared to other models, thereby explaining the faster inference. Likewise, the uncertainty-based intervention reduces the context size compared to other interventions. Running time on the ACSIncome dataset (Table 6) confirms that fairness interventions themselves are computationally inexpensive; the runtime is dominated by the choice of tabular foundation model. For practitioners, TabPFN with uncertainty-based sampling (LR) offers the best balance of fairness gains and efficiency, while TabDPT provides stronger model capacity at substantially higher cost.
| Model | Vanilla | Group Balanced | Correlation Remover | Uncertain+LR | Uncertain+TabPFN |
| TabDPT | 596.04 69.12 | 594.98 68.35 | 596.83 70.10 | 595.58 65.73 | 316.86 50.17 |
| TabICL | 69.50 0.75 | 67.84 0.85 | 70.11 1.26 | 57.05 1.24 | 96.65 1.06 |
| TabPFN | 10.71 1.27 | 10.84 1.16 | 13.94 1.35 | 8.45 0.82 | 47.18 0.65 |
| Model | Vanilla | Group Balanced | Correlation Remover | Uncertain+LR | Uncertain+TabPFN |
| TabDPT | 154.35 53.13 | 153.23 53.06 | 151.75 52.88 | 155.09 53.41 | 303.63 52.21 |
| TabICL | 45.05 0.62 | 45.23 0.04 | 44.63 0.62 | 45.09 0.29 | 45.43 1.40 |
| TabPFN | 1.85 0.87 | 1.96 1.07 | 3.22 0.32 | 1.87 0.91 | 3.65 0.76 |
Appendix D Supplementary resutls